OpenStudy (anonymous):

What is the solution of |1/2x+2| = 10

OpenStudy (anonymous):

do you have idea of solution?>

OpenStudy (anonymous):

Do you mean the answers that i am givin?

OpenStudy (anonymous):

your answer i want you to try

OpenStudy (anonymous):

i am here to help you not giving you answer to so fast

OpenStudy (anonymous):

I know. I'm just so lost and confused. Lol

OpenStudy (anonymous):

i give you hint 1=10(2x+2)

OpenStudy (anonymous):

so answer :) i correct you

OpenStudy (anonymous):

I got x=-19 over 20 ?

OpenStudy (anonymous):

perfect good job

OpenStudy (anonymous):

But that's not my answer, is it?

OpenStudy (anonymous):

replace x by the solution and tell what did you get

OpenStudy (anonymous):

1/2(-19/20+2)=

OpenStudy (anonymous):

21/40?

OpenStudy (anonymous):

no ! the answer is -19/20 like you find because if you replace it What is the solution of |1/2x+2| = 10 where is x is -19/20 it gonna give the answer 10

OpenStudy (anonymous):

understand? :)

OpenStudy (anonymous):

if you understand give me a medal ^^ if not i can explain you more :)

OpenStudy (anonymous):

So do i do 1/2(-19/20) +2 =10? Sorry...im so horrible at math. :(

OpenStudy (anonymous):

no problem we here for that you are trying to find x you find that x= -19/20 so try to do 1/(2x)+2 and put x=-19/20 if you want that 1/2(-19/20) +2= the answer 10 so it correct just try it

OpenStudy (anonymous):

Cause the answers are A -16, 24 B -24, 16 C. 16, 24 D. -16, -24. But i don't understand how to get there....

OpenStudy (anonymous):

ahh kk tell me that kk

OpenStudy (anonymous):

Okay, hold on.

OpenStudy (anonymous):

give all the given of the exercise plz

OpenStudy (anonymous):

Would it be 61/40? What i did was take 1/2 and * it by -19/20 and added 2

OpenStudy (anonymous):

just show the given and multiply choice together let me see clear

OpenStudy (anonymous):

ah wait it 1/2x + 2=10 right or 1/(2x+2)= 10 which one?